suppose you have a project that has a 0.6 chance of tripling your investment in a year and a 0.4 chance of halving your investment in a year. What is the standard deviation of the rte of return on this investment? please explain calculations
Answer #1

Let, the investment amount = $1

Expected Return = 0.6*3 + 0.4*0.5

= 2

Variance = 0.6*(3 - 2)^2 + 0.4*(0.5 - 2)^2

= 1.5

Standard Deviation = (Variance)^(1/2)

= (1.5)^(1/2)

Standard Deviation = 1.225